OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (–2 –3) and has a slope of 1/2

OpenStudy (anonymous):

Use point-slope form.\[y+3=\frac12(x+2)\\y+3=\frac12x+1\\y-\frac12x=-2\]

OpenStudy (anonymous):

\[2y-x=-4\]

OpenStudy (anonymous):

thsnks